# Hyperbolic Multiview Pretraining for Robotic Manipulation

### 05 Mar 2026 v1

@@ 3.3 Pretraining on the 3D-MOV dataset
Euclidean embeddings. For each view, we prepend a learnable ‘CLS’ token with positional encoding to aggregate global information. The resulting sequence is then processed by a stack of $N$ ViT blocks [8] (we set $N=8$), producing ‘CLS’ embeddings $\mathrm{f}^{\text{cls}}\in\mathbb{R}^{5\times 1\times D}$ and patch embeddings $\mathrm{f}^{\mathrm{p}}\in\mathbb{R}^{5\times P\times D}$. $P$ is the number of unmasked patches per view and $D$ is the embedding dimension. Mask embeddings $\mathrm{f}^{\mathrm{msk}}\in\mathbb{R}^{5\times M\times D}$ are obtained by applying a Linear layer to learnable mask tokens. $M$ is the number of masked patches per view. Unlike MAE [19], which initializes mask tokens as zero vectors [19], we randomly initialize them and apply a linear projection. It prevents mask tokens from collapsing to a single point during their mapping into hyperbolic space and ensures effective embeddings.
@@
Overall, the hyperbolic representation constraint is defined as $L_{\mathrm{hyper}}=\lambda_{\mathrm{c}}L_{\mathrm{corr}}+\lambda_{\mathrm{e1}}L_{\mathrm{etl}}(\mathrm{x}^{\text{cls}},\mathrm{x}^{\mathrm{p}})+\lambda_{\mathrm{e2}}L_{\mathrm{etl}}(\mathrm{x}^{\text{cls}},\mathrm{x}^{\mathrm{msk}})$, where the coefficients $\lambda_{\mathrm{c}}$, $\lambda_{\mathrm{e1}}$, and $\lambda_{\mathrm{e2}}$ are empirically set to $1$, $0.5$, and $0.1$, respectively.
@@
where $I_{a}^{j}$ and $\hat{I}_{i\mapsto a}^{j}$ denote the ground-truth and reconstructed pixel values of the anchor image, respectively. The overall reconstruction loss is formulated as $L_{\mathrm{recon}}=\lambda_{\mathrm{ita}}L_{\mathrm{intra}}+\lambda_{\mathrm{ite}}L_{\mathrm{inter}}$, where $\lambda_{\mathrm{ita}}$ and $\lambda_{\mathrm{ite}}$ are balancing hyperparameters, empirically set to $1$ and $0.5$, respectively.

### 12 Mar 2026 v2

@@ 3.3 Pretraining on the 3D-MOV dataset
Euclidean embeddings. For each view, we prepend a learnable ‘CLS’ token with positional encoding to aggregate global information. The resulting sequence is then processed by a stack of NN ViT blocks [8] (we set $N=8$), producing ‘CLS’ embeddings $\mathrm{f}^{\text{cls}}\in\mathbb{R}^{5\times 1\times D}$ and patch embeddings $\mathrm{f}^{\mathrm{p}}\in\mathbb{R}^{5\times P\times D}$. $P$ is the number of unmasked patches per view and $D$ is the embedding dimension. Mask embeddings $\mathrm{f}^{\mathrm{msk}}\in\mathbb{R}^{5\times M\times D}$ are obtained by applying a Linear layer to learnable mask tokens. $M$ is the number of masked patches per view. Unlike MAE [19], which initializes mask tokens as zero vectors [19], we randomly initialize them and apply a linear projection. It prevents mask tokens from collapsing to a single point during their mapping into hyperbolic space and ensures effective embeddings.
@@
Overall, the hyperbolic representation constraint is defined as $L_{\mathrm{hyper}}=\lambda_{\mathrm{c}}L_{\mathrm{corr}}+\lambda_{\mathrm{e1}}L_{\mathrm{etl}}(\mathrm{x}^{\text{cls}},\mathrm{x}^{\mathrm{p}})+\lambda_{\mathrm{e2}}L_{\mathrm{etl}}(\mathrm{x}^{\text{cls}},\mathrm{x}^{\mathrm{msk}})$, where the coefficients $\lambda_{\mathrm{c}}$, $\lambda_{\mathrm{e1}}$, and $\lambda_{\mathrm{e2}}$ are empirically set to 11, $0.5$, and $0.1$, respectively.
@@
where $I_{a}^{j}$ and $\hat{I}_{i\mapsto a}^{j}$ denote the ground-truth and reconstructed pixel values of the anchor image, respectively. The overall reconstruction loss is formulated as $L_{\mathrm{recon}}=\lambda_{\mathrm{ita}}L_{\mathrm{intra}}+\lambda_{\mathrm{ite}}L_{\mathrm{inter}}$, where $\lambda_{\mathrm{ita}}$ and $\lambda_{\mathrm{ite}}$ are balancing hyperparameters, empirically set to 11 and $0.5$, respectively.